The length of a rectangle is (x+4)cm.
The width is (x-3)cm.
The area of the rectangle is 78cm^2.
(a) Use this information to write down an equation in terms of x.
(b) Write down the length and the width of the rectangle.
I haven't been taught this yet at school but it came up in one of my practice papers :)

Answer by oscargut(2103)

a) (x+4)(x-3)=78
b)
solving: (x+4)(x-3)=78
x^2+x-12 = 78
x^2+x -90= 0
(x-9)(x+10) = 0
x = 9 or x = -10
since x must be > 0
x = 9
Length = x+4 = 13
width = x-3 = 6
You can ask me more at : mthman@gmail.com
Thanks